The function g is related to one of the parent functions g(x) = x^2 + 6 The parent function f is: f(x)= x^2 Use function notation to write g in terms of f.

1
Expert's answer
2022-06-08T01:30:46-0400

Solution:

Given parent functions "g(x)=x^2+6;" and "f(x)=x^2;"

So, variable part of "g(x)" function is equal to "f(x)" function. They are parent functions, that's why we can write "g(x)=f(x)+6;"

For example: "x=2; f(2) =2^2=4;" So, "g(2)=x^2+6=10;"

So, "g=f(2)+6=4+6=10;"

Answer:

"g=f(x)+6."